Electrical generating plants operate by generating steam at elevated temperatures and pressures then reducing the pressure in a turbine. As the pressure is reduced, high temperature steam expands (and cools) driving the turbine. The shaft work produced by the turbine is transferred to a generator to generate electricity. One limitation on steam turbines is that they cannot tolerate small amounts of water in its liquid state in the gases passing through the turbine. If the liquid content of the steam goes above some small percentage, the liquid droplets damage the turbine blades and lead to failure of the turbine.

Suppose you are operating a steam turbine where the steam leaving the turbine is at 3.99 bar (gauge). This steam contains 95 mass% vapor and 5 mass% liquid. You may assume that heat losses from the turbine are negligible for this problem.

  1. If the shaft work produced by the turbine is 1100 kJ/kg and the high pressure, high temperature steam enters the turbine at 100 bar (abs) what is the temperature of the steam entering the turbine?
  2. If water at 10ºC is supplied to the steam boiler to generate the steam that flows into the turbine, how much heat is required per kilogram of steam produced?   
Answer #1

Solution P=100 bef=lora T=1 h = ? Adialectic turbine W= 1100 KJ Kg Q0 P=3.ggbox=0.399 MPa Ol=0.95 kg Kg @ he = 604:3824k! , h#T= 643.26°C ABBox 6 Af T=10°C · Rs = 41-99 Kaling Qrequired = he-ho Qrequired (3731-7084 – 41-99) Kalen # Orequired 3689-72

In case of any doubt or you need any further help regarding this solution please comment below.

Thank you. Please upvote it.

Please............